OpenStudy (anonymous):

The length of a rectangle is twice its width. If the perimeter of the rectangle is 42 ft , find its area.

OpenStudy (rushwr):

Okai think like this The width of the rectangle is X Then the length would be 2X right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what do you guys think is next?

OpenStudy (rushwr):

then we know the perimeter of a rectangle includes 2 sides of the length and 2 sides of the width right?

OpenStudy (anonymous):

you where correct @peachpi i am sorry i got carried away i looked at a closed question and saw that you where correct

OpenStudy (rushwr):

that means the perimeter = length *2 + width *2 So 42= 2x*2 +x*2= 4x + 2x = 6x if 6x = 42 then x= 42 divided by 6 therefore X= 7 ft right?

OpenStudy (anonymous):

@peachpi i need your help

OpenStudy (rushwr):

now we know X , That is width If width = 7ft Then length = 7*2 ft= 14ft right?

OpenStudy (anonymous):

please listen to me @peachpi i am sorry

OpenStudy (rushwr):

now we know both the length and the width now its just the matter of calculating the Area Area = length * width So Area= 7ft * 14ft = ? U find that LOL :) Did u understand this ?

OpenStudy (rushwr):

@mm41 did u understand this ?

OpenStudy (anonymous):

sorta im confused on what the answer is